# Large Sets of Asymptotically Optimal and Near-Optimal Quasi-Complementary

### v1

@@ III-D The fourth construction of periodic QCSSs from Gaussian sums
$\mathbf{s}_{l}^{i,b}=\left(\mathbf{s}_{l}^{i,b}(t)\right)_{t=0}^{q-2},~{}\text{where}\ \mathbf{s}_{l}^{i,b}=\varphi_{i}(d_{l})\chi_{b}(\alpha^{t}d_{l})\chi_{1}({\alpha^{t}})~{}\text{for}\ 0\leq l\leq q-2.$
@@
$\mathcal{S}=\{\mathbf{S}^{i,b}:0\leq i\leq q-2,b\in{\mathbb{F}}_{q}\}$ (44)
@@
where $0\leq i_{1}\leq q-2,0\leq i_{2}\leq q-2,b_{1},b_{2}\in{\mathbb{F}}_{q}$. Then we consider the following cases to determine the value of $R_{\mathbf{S}^{i_{1},b_{1}},\mathbf{S}^{i_{2},b_{2}}}(\tau)$.
@@
$=$ $\left\{\begin{array}[]{ll}(q-1)\sum_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)&\text{if}~{}b_{1}-b_{2}\alpha^{\tau}=0\\ -\sum_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)&\text{if}~{}b_{1}-b_{2}\alpha^{\tau}\neq 0\end{array}\right.$
$=$ $\left\{\begin{array}[]{ll}-(q-1)&\text{if}~{}b_{1}-b_{2}\alpha^{\tau}=0,\\ 1&\text{if}~{}b_{1}-b_{2}\alpha^{\tau}\neq 0,\end{array}\right.$
@@
$=$ $\left\{\begin{array}[]{ll}\sum\limits_{y\in{\mathbb{F}}_{q}^{*}}\varphi_{i_{1}-i_{2}}(y)\sum\limits_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)&\text{if}~{}b_{1}-b_{2}\alpha^{\tau}=0\\ \begin{subarray}{c}\mbox{$\sum\limits_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)\overline{\varphi_{i_{1}-i_{2}}}((b_{1}-b_{2}\alpha^{\tau})x)\cdot$}\\ \mbox{$\sum\limits_{y\in{\mathbb{F}}_{q}^{*}}\varphi_{i_{1}-i_{2}}((b_{1}-b_{2}\alpha^{\tau})xy)\chi_{1}((b_{1}-b_{2}\alpha^{\tau})xy)$}\end{subarray}&\text{if}~{}b_{1}-b_{2}\alpha^{\tau}\neq 0\end{array}\right.$
$=$ $\left\{\begin{array}[]{ll}0&\text{if}~{}b_{1}-b_{2}\alpha^{\tau}=0,\\ \overline{\varphi_{i_{1}-i_{2}}}(b_{1}-b_{2}\alpha^{\tau})\varphi_{i_{1}-i_{2}}(1-\alpha^{\tau})G(\varphi_{i_{1}-i_{2}},\chi_{1})G(\overline{\varphi_{i_{1}-i_{2}}},\chi_{1})&\text{if}~{}b_{1}-b_{2}\alpha^{\tau}\neq 0,\end{array}\right.$
by the orthogonal relation of the additive and multiplicative characters and the definition of Gaussian sum. It is known that $G(\overline{\varphi_{i_{1}-i_{2}}},\chi_{1})=\varphi_{i_{1}-i_{2}}(-1)\overline{G(\varphi_{i_{1}-i_{2}},\chi_{1})}$. Then
@@
$\vartheta_{opt}=\sqrt{\frac{(q-1)^{5}}{q(q-1)^{2}-1}}.$
@@
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{max}}{\vartheta_{opt}}=\lim_{q\rightarrow+\infty}\frac{q}{\sqrt{\frac{(q-1)^{5}}{q(q-1)^{2}-1}}}=1.$
@@ IV Concluding remarks
In this paper, we presented five new families of asymptotically optimal or near-optimal periodic QCSSs with large set sizes from some special polynomials and Gaussian sums. Specially, we constructed the first two families of periodic QCSSs with set size $\Theta(q^{3})$ and flock size $\Theta(q)$, wher $q$ is a power of a prime. Most of the QCSSs constructed in this paper have very small alphabet size. The parameters of the QCSSs in this paper improve those of known ones in [10, 22]. From the figures in this paper, some complementary sequences have zero nontrivial autocorrelation magnitude. These advantages indicate that our QCSSs have potential applications in MC-CDMA communication systems.

### v2

@@ III-D The fourth construction of periodic QCSSs from Gaussian sums
$\mathbf{s}_{l}^{i,b}=\left(\mathbf{s}_{l}^{i,b}(t)\right)_{t=0}^{q-2},\leavevmode\nobreak\ \text{where}\ \mathbf{s}_{l}^{i,b}=\varphi_{i}(d_{l})\chi_{b}(\alpha^{t}d_{l})\chi_{1}({\alpha^{t}})\leavevmode\nobreak\ \text{for}\ 0\leq l\leq q-2.$
@@
$\mathcal{S}=\left\{\mathbf{S}^{i,b}:0\leq i\leq q-2,b\in{\mathbb{F}}_{q}\right\}$ (44)
@@
where $0\leq i_{1}\leq q-2,0\leq i_{2}\leq q-2,b_{1},b_{2}\in{\mathbb{F}}_{q}$. Then we consider the following cases to determine the value distribution of $R_{\mathbf{S}^{i_{1},b_{1}},\mathbf{S}^{i_{2},b_{2}}}(\tau)$.
@@
$=$ $\left\{\begin{array}[]{ll}(q-1)\sum_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}=0\\ -\sum_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}\neq 0\end{array}\right.$
$=$ $\left\{\begin{array}[]{ll}-(q-1)&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}=0,\\ 1&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}\neq 0,\end{array}\right.$
@@
$=$ $\left\{\begin{array}[]{ll}\sum\limits_{y\in{\mathbb{F}}_{q}^{*}}\varphi_{i_{1}-i_{2}}(y)\sum\limits_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}=0\\ \begin{subarray}{c}\mbox{$\sum\limits_{x\in{\mathbb{F}}_{q}^{*}}\chi_{1}((1-\alpha^{\tau})x)\overline{\varphi_{i_{1}-i_{2}}}((b_{1}-b_{2}\alpha^{\tau})x)\cdot$}\\ \mbox{$\sum\limits_{y\in{\mathbb{F}}_{q}^{*}}\varphi_{i_{1}-i_{2}}((b_{1}-b_{2}\alpha^{\tau})xy)\chi_{1}((b_{1}-b_{2}\alpha^{\tau})xy)$}\end{subarray}&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}\neq 0\end{array}\right.$
$=$ $\left\{\begin{array}[]{ll}0&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}=0,\\ \overline{\varphi_{i_{1}-i_{2}}}(b_{1}-b_{2}\alpha^{\tau})\varphi_{i_{1}-i_{2}}(1-\alpha^{\tau})G(\varphi_{i_{1}-i_{2}},\chi_{1})G(\overline{\varphi_{i_{1}-i_{2}}},\chi_{1})&\text{if}\leavevmode\nobreak\ b_{1}-b_{2}\alpha^{\tau}\neq 0,\end{array}\right.$
by the orthogonal relation of the additive and multiplicative characters and the definition of Gaussian sums. It is known that $G(\overline{\varphi_{i_{1}-i_{2}}},\chi_{1})=\varphi_{i_{1}-i_{2}}(-1)\overline{G(\varphi_{i_{1}-i_{2}},\chi_{1})}$. Then
@@
$\vartheta_{{\text{opt}}}=\sqrt{\frac{(q-1)^{5}}{q(q-1)^{2}-1}}.$
@@
$\lim_{q\rightarrow+\infty}\frac{\vartheta_{\max}}{\vartheta_{{\text{opt}}}}=\lim_{q\rightarrow+\infty}\frac{q}{\sqrt{\frac{(q-1)^{5}}{q(q-1)^{2}-1}}}=1.$
@@ IV Concluding remarks
In this paper, we presented five new families of asymptotically optimal or near-optimal periodic QCSSs with large set sizes from some special polynomials and Gaussian sums. Specially, we constructed the first two families of periodic QCSSs with set size $\Theta(K^{3})$ and flock size $K$. Most of the QCSSs constructed in this paper have very small alphabet size. The parameters of the QCSSs in this paper improve those of known ones in [10, 22]. From the figures in this paper, some complementary sequences have zero nontrivial autocorrelation magnitude. These advantages indicate that our QCSSs have potential applications in MC-CDMA communication systems.
@@
Declarations
Conflict of interest The authors declare that they have no conflicts of interest relevant to the content of this article.